OpenStudy (anonymous):

Matt is constructing two similar triangles for an art project. The drawing below shows Matt's plans, but there is an error in his drawing. What changes would he make to the dimensions to change the error? Explain your reasoning using complete sentences

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

i'm doing

OpenStudy (alexandervonhumboldt2):

13/2.5=5.2 43/8.4=5.1190476190476190476190476190476 20.8/4=5.2 Similar triangles have proportonla sides. proportions are equal. So demensons of DE or AC are wrong

OpenStudy (anonymous):

Thanks you so much, your amazing!!!!

OpenStudy (alexandervonhumboldt2):

sorry but dont know how to find which side is wrong